# Matching Hadronization and Perturbative Evolution: The Cluster Model in Light of Infrared Shower Cutoff Dependence

### v1

@@ 1 Introduction
Multi-purpose Monte Carlo event generators (MCs) are indispensable tools to describe realistic, fully detailed hadronic final states for essentially all processes at collider experiments. Their underlying structure and components reflect the large hierarchies of energy scales involved in these processes. These scale hierarchies are also the basis of numerous factorization theorems in analytic QCD approaches, which are partly also reflected in the MC components. After the determination of hard scattering cross sections using matrix elements obtained from fixed-order perturbation theory, all-order leading contributions coming from the radiation of soft and collinear partons are resummed by parton shower algorithms. The latter evolve the incoming and outgoing energetic partons participating in the hard scattering to lower scales tied to an ordering variable. This evolution terminates when the ordering variable reaches a cut-off value, which we refer to as $Q_{0}$, and which defines a low scale kinematic restriction on the gluon radiation and gluon branching into quarks. The value of $Q_{0}$ is typically in the range between $0.5$ and $2$ GeV. Phenomenological models of hadronization then describe how hadrons are formed out of the partonic final states that have emerged from the parton shower. The parameters of these hadronization models are fixed through the tuning procedure which is based on a fit to a reference data set. In current state-of-the-art MCs it is common that the shower cutoff $Q_{0}$ is also treated as a hadronization parameter. This means that the value of $Q_{0}$ is fixed by demanding best agreement with the reference data. In collisions of extended objects, such as protons, also multi-parton interactions taking place in the initial stages of the collision are simulated. The modelling of multi-parton interactions is, however, not subject of the present paper.
@@
One of the major recent advances in the development of MCs has been to include fixed-order (QCD) corrections for the description of the hard scattering process and the production of additional hard jets through NLO matching and multijet merging algorithms. These developments need to be complemented by more accurate parton shower algorithms to improve the description of the resummed higher order corrections arising from the soft and collinear dynamics. Such improvements for a large class of observables have only recently received more attention: While, since a long time, parton shower approaches such as coherent branching can be rigorously (i.e. analytically and numerically) proven to be accurate at the next-to-leading logarithmic (NLL) level for dijet-based $e^{+}e^{-}$ event-shape variables such as thrust or other event-shapes related to jet masses in the peak region Catani:1992ua ; Hoang:2018zrp ; Bewick:2019rbu ; Bewick:2021nhc , similar progress for dipole parton showers, which are more convenient for matching and merging, has only recently been achieved, see e.g. Dasgupta:2020fwr ; Forshaw:2020wrq ; Herren:2022jej for first steps in this direction. Going even beyond this level of accuracy for the parton shower evolution is also subject to an active area of research, but the implementation of these developments in the framework of multi-purpose MCs applicable for experimental analyses may likely still require significant time.
@@
Inspired by earlier studies of some of the authors in Ref. Proceedings:2018jsb , we argue that accurate and improved parton shower algorithms to be employed for MC predictions should not be considered independently of the MC hadronization models. In principle this is already obvious at the purely practical level in the context of the tuning procedure of an event generator to a set of reference data, since the tuning traditionally considers a combined fit of parton shower and hadronization model parameters. However, what we mean is, that eventually one should go beyond this practical level by demanding in addition that the combined matrix-element and parton shower partonic description, on one side, and the effects of the hadronization model on the other, by themselves have a well-defined field theoretic meaning in the context of QCD with a systematic and controllable scheme-dependence. It is obvious that this is particularly important when combining hadronization corrections extracted from MCs with high precision perturbative QCD predictions for the determination of hadron level eventshape distribution, or for the interpretation of $m_{t}^{\rm MC}$ obtained from the direct measurements. This implies that (i) the field theoretic meaning of the parton level description and the hadronization effects should not depend on any parameter subject to tuning hadronization effects and (ii) in particular that the parton shower infrared cutoff $Q_{0}$ should be considered as a factorization scale. This view has also been advocated in Ref. Platzer:2022jny , where the same reasoning has been applied for the construction of an algorithm for parton branching at the amplitude level Forshaw:2019ver based on an underlying factorization of jet cross sections. The way how the shower cutoff is implemented defines a particular scheme of this factorization, and the factorization property ensures that the hadron level description is scheme-invariant. This in particular implies that, at least within some limited $Q_{0}$ interval, the combined partonic and hadron-model prediction for infrared sensitive quantities should be independent of the $Q_{0}$ value. As it is probably too ambitious to demand exact $Q_{0}$-independence, in practice at least a systematic cancellation between the dominant linear $Q_{0}$-dependencies of the parton level predictions and of the associated hadronization corrections (at the precision level of the parton shower) should be realized. This $Q_{0}$-insensitivity with respect to linear $Q_{0}$ contributions is what we refer to as ’$Q_{0}$-independence’ in this article.
@@
Within such a framework the analytic properties of a (NLO matched and) NLL accurate parton shower can be transferred in a fully controllable way to the hadron level description provided by the MC generator and new avenues to scrutinize the combined action of parton level and hadron level descriptions provided by the MCs are made possible. This level of control is also mandatory to study the impact of the shower cutoff $Q_{0}$ dependent top quark mass parameter $m_{Q}^{\rm CB}(Q_{0})$, which was recently proven to emerge from using the coherent branching algorithm for massive quark initiated $e^{+}e^{-}$ event-shape distributions Hoang:2018zrp , in the context of MC simulations for boosted top pair production.
@@
In this article we further promote the idea of the shower cutoff $Q_{0}$ being a factorization scale by providing an actual implementation of a hadronization model acting in this direction. This entails that the (linear) evolution in $Q_{0}$ of the parton level description and the corresponding hadronization effects individually follow the concrete predictions of QCD perturbation theory, yielding $Q_{0}$-independent hadron level descriptions in a controllable and systematic manner. To be definite, we carry out our considerations in the context of the angular ordered parton shower and cluster hadronization model as implemented in the Herwig 7 MC generator. Our new model will become available with an upcoming Herwig 7.4 release. We focus in particular on the $e^{+}e^{-}$ dijet event-shape distribution 2-jettiness. For the latter it is known since a long time ago Catani:1992ua ; Contopanagos:1996nh ; Almeida:2014uva that the coherent branching algorithm, which is the basis of the angular ordered Herwig parton shower, is NLL precise. For the 2-jettiness distribution the dominant linear shower cut $Q_{0}$-dependence of the parton shower description in the dijet limit can also be determined analytically at NLO (${\cal O}(\alpha_{s})$) and the Herwig 7.2 parton shower has been shown to be fully consistent with these analytic results also for quasi-collinear massive quarks in Ref. Hoang:2018zrp . Using analytic results for this linear $Q_{0}$-dependence and the association of the generator’s parton-to-hadron level migration matrix $T$ with the dijet shape function $S_{\rm had}$ appearing in QCD factorization and soft-collinear effective theory (SCET) for 2-jettiness in the dijet region Lee:2006nr , we can derive a QCD constraint concerning the migration matrix $T$, which can then be tested in detail. This constraint involves the linear $Q_{0}$-dependence of the first moment of the transfer matrix and demands in addition independence concerning the c.m. energy $Q$.
@@
We find that the default cluster hadronization model of Herwig 7.2 does not satisfy this QCD constraint in a consistent manner.111 In this respect, Herwig 7.3, which has been released recently, behaves similar to Herwig 7.2. This motivates the construction of a novel and improved cluster hadronization model, which we call the “dynamical cluster model”. The dynamical cluster model satisfies the QCD constraints more accurately and leads to a significantly improved shower cut $Q_{0}$-independence of the hadron level description in the physically important shower cutoff interval $21~{}\mbox{GeV}<Q_{0}<$ 2 GeV. While even the behavior of the new dynamic hadronization model with respect to $Q_{0}$-variations is not yet perfect, the results of this article provide first important steps towards giving the hadronization effects provided by MC simulations a well-defined QCD meaning with a controllable scheme-dependence, and a stepping stone to achieve hadronization models fully consistent with QCD factorization, as outlined in Platzer:2022jny . Eventually variations of the shower cut may also be used as an instrument to quantity the theoretical uncertainty of event generator predictions in analogy to the canonical renormalization or factorization scale variations in analytic calculations. In this article we focus on hadronization in the presence of light quarks, which is a vital step towards understanding the interplay of the infrared cutoff, hadronization and the shower cutoff dependence of heavy quark mass parameters in an upcoming paper ourMCtopmasspaper .
@@ 2.1 General remarks
Coherent branching is a parton shower algorithm which is based on $1\to 2$ parton splitting processes incorporating QCD coherence through the ordering in an emission angular variable $\tilde{q}$, see e.g. the classic articles Catani:1990rr ; Catani:1992ua . In this work we use the implementation of coherent branching in the Herwig event generator, as previously described and analyzed in Gieseke:2003rz ; Hoang:2018zrp . For global observables in simple 2-jet processes such as event-shapes distributions in $e^{+}e^{-}\to$ hadrons, coherent branching is NLL precise. Here we do not discuss details of Herwig’s coherent branching algorithm [^2] We refer the reader to Ref. Hoang:2018zrp for a detailed discussion of features in Herwig’s coherent branching parton sbower relevant for the shower cutoff analyses in this article. but focus on its main features relevant for the kinematic properties of the emerging partonic final state and its characteristics relevant for the onset of the simulated hadronization dynamics taking place after the parton shower has terminated.
The first important aspect related to hadronization is that the coherent branching algorithm generates colour structures which are largely compatible with the space-time structure of hadronization in the sense that large-angle soft gluons will be emitted first in the parton shower evolution. Upon hadronization this effectively isolates the colors of the outgoing hard partons in the form of jets at the expense of forming a few, soft hadrons transverse to the hard jet momenta. The latter feature is also the basis of the analytic factorization methods for large-angle soft and collinear radiation dynamics such as in Soft-Collinear Effective Theory (SCET) Bauer:2002nz . The second important aspect is the preconfining property of the coherent QCD evolution. This property requires that, in the large-$N_{c}$ limit, color singlets (i.e. colour connected quark-antiquark pairs which form at the end of the shower through additional $g\to q\bar{q}$ branchings to be discussed in more detail in Sec. 4 and 5) acquire a universal invariant mass spectrum that is peaked at scales similar to the infrared cutoff $Q_{0}$ of the shower evolution and falls off exponentially for larger invariant masses. These ’clusters’, expressing the fact that color correlations are very local in phase space, can be interpreted as excited hadronic states and form the basic degree of freedom of the cluster hadronization model.
@@ 2.2 Kinematic Reconstruction, Reshuffling and True Parton Level
The (on-shell) masses of the final state partons emerging from the kinematic reconstruction procedure are in principle free parameters. In the context of processes involving only massless quarks and gluons, one would expect them to be zero. This is what we refer to as the true parton level, which can also be related to massless parton final states in standard computations in perturbative QCD. In general the true parton level is associated to the quarks having current quark masses $\hat{m}_{i}$. [^3] This true final state parton level has been employed and analyzed in detail already in our previous work Hoang:2018zrp . In practice, for a lght quarks a mass of $10$ keV is adopted. In that previous work we have also analyzed in depth how the parton shower algorithm affects the interpretation of the mass parameter for heavy quarks in relation to mass renormalization schemes. However, the initial steps to interface the parton level to the cluster hadronization model requires (much larger) constituent quark masses $m_{i}$ for all final state quarks. These constituent quark masses are parameters of the hadronization model and are constrained such that a cluster is kinematically allowed to decay at least into a pair of the lightest hadrons. Furthermore, the cluster hadronization model involves a branching of each gluon into $q\bar{q}$ pairs as the basis to form the initial clusters. The model thus assigns the gluons a mass $m_{g}$, such that this $g\to q\bar{q}$ splitting process is kinematically allowed at least for the lightest quarks with respect to their constituent masses. [^4] Below we also refer to the gluon mass $m_{g}$ as a constituent mass for simplicity.
@@ 3.1 General Formulation
Here ${\rm d}\phi_{m}(p_{1},...,p_{m}|Q)$ is the integration measure of the final-state hadron momenta and the term ${\rm d}\sigma_{P,n,c}(q_{1},..,q_{n}|Q;Q_{0})$ stands for the partonic cross section, including the partonic phase space integration of total momentum $Q$ and a certain colour-flow $c$. The term $S_{mn,c}(p_{1},...,p_{m}|q_{1},...,q_{n};Q_{0})$ represents the action of the hadronization model in converting $n$ partons into $m$ hadrons subject to a given model and momentum mapping inherent to it. $W(p_{1}...,p_{m})$ in turn is the observable’s definition at hadron level. In practice for an event generator we have $m>n$. We also stress the fact that such a probabilistic factorization is merely possible in presence of the large-$N$ limit and more generally would involve the presence of a colour flow in the parton level amplitude and its conjugate, see Platzer:2022jny for more details. In the above form it is clear that the demand of $Q_{0}$-independence implies sets of evolution equations for both the partonic and hadronic factor, which mix different partonic multiplicities, a generic feature which a hadronization model consistent with $Q_{0}$ independence must respect. To this end, we can use the evolution of the parton shower, which we can schematically write for a gluon exchange or emission with momentum $q_{n}$ in terms of a virtual $V_{c}(q_{1},..,q_{n})$ and real emission contributon $R_{c,c^{\prime}}(q_{1},..,q_{n})$ as
$\frac{\partial}{\partial Q_{0}}{\rm d}\sigma_{P,n,c}(q_{1},...,q_{n}|Q;Q_{0})=\int\left(V_{c}(q_{1},..,q_{n}){\rm d}\sigma_{P,n,c}(q_{1},...,q_{n}|Q;Q_{0})\right.\\ \left.-\sum_{c^{\prime}}R_{c,c^{\prime}}(q_{1},..,q_{n}){\rm d}\sigma_{P,n-1,c^{\prime}}(q_{1},...,q_{n-1}|Q;Q_{0})\right){\rm d}^{d}q_{n}\ ,$ (2)
@@
where $V$ and $R$ include the definition of some momentum mapping, and can eventually be expressed in terms of splitting functions. Demanding that the hadronic cross section is independent of $Q_{0}$, one can then derive an evolution equation for $S_{mn}$ Platzer:2022jny . For the present purpose we note that Eq. (1) is an accurate analytic model of an event generator, and Eq. (2) one of parton shower evolution. We can also ask the question how such a factorization then gives rise to observable-specific hadronization corrections, which can be extracted in a MC sense by snapshots of events at the parton level and at the level of the final hadronic states. For $e^{+}e^{-}$ event-shapes this is discussed in the next subsection. Furthermore, we are naturally led to conclude that $Q_{0}$ should not be a tuning parameter and that we should investigate the $Q_{0}$ dependence of $S_{mn}$, for example, by studying the dependence of the tuned hadronization model parameters on variations of $Q_{0}$. As already mentioned in the introduction, the least we need to demand from an improved $S_{mn}$ is a smooth change across multiplicities in the sense that higher partonic multiplicities at lower $Q_{0}$ lead to comparable effect as smaller multiplicities at higher $Q_{0}$ do: the dynamics governing $S_{mn}$ near the infrared cutoff need to be the same as that of the shower close to the infrared cutoff. This puts severe constraints on the onset of the hadronization model, to be discussed in Sec. 4 and Sec. 5. Using the evolution equations implied for $S_{mn}$ for the construction of an evolving model is subject of additional ongoing work.

### v2

@@ 1 Introduction
Multi-purpose Monte Carlo event generators (MCs) are indispensable tools to describe realistic, fully detailed hadronic final states for essentially all processes at collider experiments. Their underlying structure and components reflect the large hierarchies of energy scales involved in these processes. These scale hierarchies are also the basis of numerous factorization theorems in analytic QCD approaches, which are partly also reflected in the MC components. After the determination of hard scattering cross sections using matrix elements obtained from fixed-order perturbation theory, all-order leading contributions coming from the radiation of soft and collinear partons are resummed by parton shower algorithms. The latter evolve the incoming and outgoing energetic partons participating in the hard scattering to lower scales tied to an ordering variable. This evolution terminates when the ordering variable reaches a cut-off value, which we refer to as $Q_{0}$, and which defines a low scale kinematic restriction on the gluon radiation and gluon branching into quarks and gluons. The value of $Q_{0}$ is typically in the range between $0.5$ and $2$ GeV. Phenomenological models of hadronization then describe how hadrons are formed out of the partonic final states that have emerged from the parton shower. The parameters of these hadronization models are fixed through the tuning procedure which is based on a fit to a reference data set. In current state-of-the-art MCs it is common that the shower cutoff $Q_{0}$ is also treated as a hadronization parameter. This means that the value of $Q_{0}$ is fixed by demanding best agreement with the reference data. In collisions of extended objects, such as protons, also multi-parton interactions taking place in the initial stages of the collision are simulated. The modelling of multi-parton interactions is, however, not subject of the present paper.
@@
One of the major recent advances in the development of MCs has been to include fixed-order (QCD) corrections for the description of the hard scattering process and the production of additional hard jets through NLO matching and multijet merging algorithms. These developments need to be complemented by more accurate parton shower algorithms to improve the description of the resummed higher order corrections arising from the soft and collinear dynamics. Such improvements for a large class of observables have only recently received more attention: While, since a long time, parton shower approaches such as coherent branching can be rigorously (i.e. analytically and numerically) proven to be accurate at the next-to-leading logarithmic (NLL) level for dijet-based $e^{+}e^{-}$ event-shape variables such as thrust or other event-shapes related to jet masses in the peak region Catani:1992ua ; Hoang:2018zrp ; Bewick:2019rbu ; Bewick:2021nhc , similar progress for dipole parton showers, which are more convenient for matching and merging, has only recently been achieved, see e.g. Dasgupta:2020fwr ; Forshaw:2020wrq ; Nagy:2020rmk ; Herren:2022jej ; vanBeekveld:2022zhl ; Hoche:2024dee for recent proposals and implementations. Going even beyond this level of accuracy for the parton shower evolution is also subject to an active area of research, but the implementation of these developments in the framework of multi-purpose MCs applicable for experimental analyses may likely still require significant time. [^1] A first study vanBeekveld:2024wws has appeared just after the first version of this work has been submitted.
@@
Inspired by earlier studies of some of the authors in Ref. Proceedings:2018jsb , we argue that accurate and improved parton shower algorithms to be employed for MC predictions should not be considered independently of the MC hadronization models. In principle this is already obvious at the purely practical level in the context of the tuning procedure of an event generator to a set of reference data, since the tuning traditionally considers a combined fit of parton shower and hadronization model parameters. However, what we mean is, that eventually one should go beyond this practical level by demanding in addition that the combined matrix-element and parton shower partonic description, on one side, and the effects of the hadronization model on the other, by themselves have a well-defined field theoretic meaning in the context of QCD with a systematic and controllable scheme-dependence. It is obvious that this is particularly important when combining hadronization corrections extracted from MCs with high precision perturbative QCD predictions for the determination of hadron level eventshape distribution, or for the interpretation of $m_{t}^{\rm MC}$ obtained from the direct measurements. This implies that (i) the field theoretic meaning of the parton level description and the hadronization effects should not depend on any parameter subject to tuning hadronization effects and (ii) in particular that the parton shower infrared cutoff $Q_{0}$ should be considered as a factorization scale.
This view has also been advocated in Ref. Platzer:2022jny , where the same reasoning has been applied for the construction of an algorithm for parton branching at the amplitude level Forshaw:2019ver based on an underlying factorization of jet cross sections. The way how the shower cutoff is implemented (in terms of an infrared restriction related to an evolution variable) defines a particular scheme of this factorization, and the factorization property ensures that the hadron level description is scheme-invariant. This in particular implies that, at least within some limited $Q_{0}$ interval, the combined partonic and hadron-model prediction for infrared sensitive quantities should be independent of the $Q_{0}$ value and the hadronization model should be capable to describe the perturbative evolution in $Q_{0}$. The essential merrit of a model satisfying this property would be that its genuine hadronization effects related to $\Lambda_{\rm QCD}$ would become universal and thus well-defined within QCD. In contrast a model not having the property (or a model where this property has never been examined in a dedicated way), can yield hadronization effects that are not consistent with QCD. As it is probably too ambitious to demand exact $Q_{0}$-independence, in practice at least a sizeable systematic cancellation between the dominant linear $Q_{0}$-dependencies of the parton level predictions and of the associated hadronization corrections (at the precision level of the parton shower) should be realized. This level of $Q_{0}$-insensitivity with respect to linear $Q_{0}$ contributions is what we refer to as ’$Q_{0}$-independence’ in this article.
@@
Within such a framework the analytic properties of a (NLO matched and) NLL accurate parton shower could be transferred in a controllable way to the hadron level description provided by the MC generator and new avenues to scrutinize the combined action of parton level and hadron level descriptions provided by the MCs would be made possible. This level of control is also mandatory to study the impact of the shower cutoff $Q_{0}$ dependent top quark mass parameter $m_{Q}^{\rm CB}(Q_{0})$, which was recently proven to emerge from using the coherent branching algorithm for massive quark initiated $e^{+}e^{-}$ event-shape distributions Hoang:2018zrp , in the context of MC simulations for boosted top pair production.
@@
In this article we further promote the idea of the shower cutoff $Q_{0}$ being a factorization scale by providing an actual implementation of a hadronization model acting in this direction. This entails that the (linear) evolution in $Q_{0}$ of the parton level description and the corresponding hadronization effects individually follow the concrete predictions of QCD perturbation theory with good precision, yielding hadron level descriptions that are less $Q_{0}$-dependent in a controllable manner. To be definite, we carry out our considerations in the context of the angular ordered parton shower and cluster hadronization model as implemented in the Herwig 7 MC generator Bellm:2015jjp ; Bellm:2019zci ; Bewick:2023tfi . Our new model will become available with an upcoming Herwig 7.4 release. We focus in particular on the $e^{+}e^{-}$ dijet event-shape distribution 2-jettiness. For the latter it is known since a long time ago Catani:1992ua ; Contopanagos:1996nh ; Almeida:2014uva that the coherent branching algorithm, which is the basis of the angular ordered Herwig parton shower, is NLL precise. For the 2-jettiness distribution the dominant linear shower cut $Q_{0}$-dependence of the angular ordered parton shower description in the dijet limit has also been determined analytically at NLO (${\cal O}(\alpha_{s})$), and the Herwig 7.2 parton shower has been shown to be fully consistent with these analytic results also for quasi-collinear massive quarks in Ref. Hoang:2018zrp .222 The results obtained in Ref. Hoang:2018zrp were possible due to the global character of jet mass based event-shapes such as 2-jettiness (or thrust), and because of the knowledge on their factorization at the hadron level. At this time results along the same lines have not been determined for other types of observables. Using the analytic results for this linear $Q_{0}$-dependence and the association of the generator’s parton-to-hadron level migration matrix $T$ with the dijet shape function $S_{\rm had}$ appearing in QCD factorization and soft-collinear effective theory (SCET) for 2-jettiness in the dijet region Lee:2006nr , we can derive a QCD constraint concerning the migration matrix $T$, which can then be tested in detail. This constraint involves the linear $Q_{0}$-dependence of the first moment of the transfer matrix and demands in addition independence concerning the c.m. energy $Q$.
@@
We find that the default cluster hadronization model of Herwig 7.2 does not satisfy this QCD constraint in a consistent manner.333 In this respect, Herwig 7.3 Bewick:2023tfi , which has been released recently, behaves similar to Herwig 7.2 Bellm:2019zci . This motivates the construction of a novel and improved cluster hadronization model, which we call the “dynamical cluster model”. The dynamical cluster model satisfies the QCD constraints more accurately and leads to a significantly improved shower cut $Q_{0}$-independence of the hadron level description in the physically important shower cutoff interval $21~{}\mbox{GeV}<Q_{0}<$ 2 GeV. While even the behavior of the new dynamic hadronization model with respect to $Q_{0}$-variations is not yet perfect, the results of this article provide first important steps towards giving the hadronization effects provided by MC simulations a well-defined QCD meaning with a controllable scheme-dependence, and a stepping stone to achieve hadronization models fully consistent with QCD factorization, as outlined in Platzer:2022jny . Eventually variations of the shower cut may also be used as an instrument to quantify the theoretical uncertainty of event generator predictions in analogy to the canonical renormalization or factorization scale variations in analytic calculations. In this article we focus on hadronization in the presence of light quarks, which is a vital step towards understanding the interplay of the infrared cutoff, hadronization and the shower cutoff dependence of heavy quark mass parameters in an upcoming paper ourMCtopmasspaper .
@@ 2.1 General remarks
Coherent branching is a parton shower algorithm which is based on $1\to 2$ parton splitting processes incorporating QCD coherence through the ordering in an emission angular variable $\tilde{q}$, see e.g. the classic articles Catani:1990rr ; Catani:1992ua . In this work we use the implementation of coherent branching in the Herwig event generator, as previously described and analyzed in Gieseke:2003rz ; Hoang:2018zrp . For global observables in simple 2-jet processes such as event-shapes distributions in $e^{+}e^{-}\to$ hadrons, coherent branching is NLL precise. Here we do not discuss details of Herwig’s coherent branching algorithm [^4] We refer the reader to Ref. Hoang:2018zrp for a detailed discussion of features in Herwig’s coherent branching parton shower relevant for the shower cutoff analyses in this article. but focus on its main features relevant for the kinematic properties of the emerging partonic final state and its characteristics relevant for the onset of the simulated hadronization dynamics taking place after the parton shower has terminated.
The first important aspect related to hadronization is that the coherent branching algorithm generates colour structures which are largely compatible with the space-time structure of hadronization in the sense that large-angle soft gluons will be emitted first in the parton shower evolution. Upon hadronization this effectively isolates the colors of the outgoing hard partons in the form of jets at the expense of forming a few, soft hadrons transverse to the hard jet momenta. The latter feature is also the basis of the analytic factorization methods for large-angle soft and collinear radiation dynamics such as in Soft-Collinear Effective Theory (SCET) Bauer:2002nz . The second important aspect is the preconfining property of the coherent QCD evolution. This property requires that, in the large-$N_{c}$ limit, color singlets (i.e. colour connected quark-antiquark pairs which form at the end of the shower through additional $g\to q\bar{q}$ branchings to be discussed in more detail in Sec. 4 and 5) acquire a universal invariant mass spectrum that is peaked at scales similar to the infrared cutoff $Q_{0}$ of the shower evolution and falls off exponentially for larger invariant masses. These ’clusters’, expressing the fact that color correlations are very local in phase space, can be interpreted as excited hadronic states and form the basic degree of freedom of the cluster hadronization model. The way how the coherent evolution builds up colour structure in the large-$N_{c}$ limit also reflect the time scales involed in the build-up of jets and the way colour charges of the inital $q\bar{q}$ pair are isolated from each other, predicting few soft hadrons transverse to the hard jet directions. Coherent branching and the cluster model, together with global observables which can reliably be predicted using coherence, thus fit very well to study the interface between shower and hadronization. [^5] Non-global effects might not be predicted in an entirely reliable way from coherent branching, however hadronization dynamics in this case are also far from well-understood.
@@ 2.2 Kinematic Reconstruction, Reshuffling and True Parton Level
The (on-shell) masses of the final state partons emerging from the kinematic reconstruction procedure are in principle free parameters. In the context of processes involving only massless quarks and gluons, one would expect them to be zero. This is what we refer to as the true parton level, which can also be related to massless parton final states in standard computations in perturbative QCD. In general the true parton level is associated to the quarks having current quark masses $\hat{m}_{i}$. [^6] This true final state parton level has been employed and analyzed in detail already in our previous work Hoang:2018zrp . In practice, for a light quarks a mass of $10$ keV is adopted. In that previous work we have also analyzed in depth how the parton shower algorithm affects the interpretation of the mass parameter for heavy quarks in relation to mass renormalization schemes. However, the initial steps to interface the parton level to the cluster hadronization model requires (much larger) constituent quark masses $m_{i}$ for all final state quarks. These constituent quark masses are parameters of the hadronization model and are constrained such that a cluster is kinematically allowed to decay at least into a pair of the lightest hadrons. Furthermore, the cluster hadronization model involves a branching of each gluon into $q\bar{q}$ pairs as the basis to form the initial clusters. The model thus assigns the gluons a mass $m_{g}$, such that this $g\to q\bar{q}$ splitting process is kinematically allowed at least for the lightest quarks with respect to their constituent masses. [^7] Below we also refer to the gluon mass $m_{g}$ as a constituent mass for simplicity.
@@ 3.1 General Formulation
Here ${\rm d}\phi_{m}(p_{1},...,p_{m}|Q)$ is the integration measure of the final-state hadron momenta and the term ${\rm d}\sigma_{P,n,c}(q_{1},..,q_{n}|Q;Q_{0})$ stands for the partonic cross section, including the partonic phase space integration of total momentum $Q$ and a certain colour-flow $c$. The term $S_{mn,c}(p_{1},...,p_{m}|q_{1},...,q_{n};Q_{0})$ represents the action of the hadronization model in converting $n$ partons into $m$ hadrons subject to a given model and momentum mapping inherent to it. $W(p_{1}...,p_{m})$ in turn is the observable’s definition at hadron level. In practice for an event generator we have $m>n$. We also stress the fact that such a probabilistic factorization is merely possible in presence of the large-$N$ limit and more generally would involve the presence of a colour flow in the parton level amplitude and its conjugate, see Platzer:2022jny for more details. In the above form it is clear that the demand of $Q_{0}$-independence implies sets of evolution equations for both the partonic and hadronic factor, which mix different partonic multiplicities, a generic feature which a hadronization model consistent with $Q_{0}$ independence must respect. To this end, we can use the evolution of the parton shower, which we can schematically write for a gluon exchange or emission with momentum $q_{n}$ in terms of a virtual $V_{c}(q_{1},..,q_{n};Q_{0})$ and real emission contributon $R_{c,c^{\prime}}(q_{1},..,q_{n}|p_{1},...,p_{n-1};Q_{0})$ as
$Q_{0}\frac{\partial}{\partial Q_{0}}{\rm d}\sigma_{P,n,c}(q_{1},...,q_{n}|Q;Q_{0})=V_{c}(q_{1},..,q_{n};Q_{0}){\rm d}\sigma_{P,n,c}(q_{1},...,q_{n}|Q;Q_{0})\\ -\sum_{c^{\prime}}\int R_{c,c^{\prime}}(q_{1},..,q_{n}|p_{1},...,p_{n-1};Q_{0}){\rm d}\sigma_{P,n-1,c^{\prime}}(p_{1},...,p_{n-1}|Q;Q_{0}){\rm d}^{d}p_{1}\cdots{\rm d}^{d}p_{n}\ ,$ (2)
@@
where $V$ and $R$ include the definition of some momentum mapping, and can eventually be expressed in terms of splitting functions. Demanding that the hadronic cross section is independent of $Q_{0}$, one can then derive an evolution equation for $S_{mn}$ Platzer:2022jny . This evolution demands that the hadronization model should be able to mirror the dynamics of the parton shower in evolving towards a low-scale process of producing hadrons. For the present purpose we note that Eq. (1) is an accurate analytic model of an event generator, and Eq. (2) one of parton shower evolution. We can also ask the question how such a factorization then gives rise to observable-specific hadronization corrections, which can be extracted in a MC sense by snapshots of events at the parton level and at the level of the final hadronic states. For $e^{+}e^{-}$ event-shapes this is discussed in the next subsection. Furthermore, we are naturally led to conclude that $Q_{0}$ should not be a tuning parameter and that we should investigate the $Q_{0}$ dependence of $S_{mn}$, for example, by studying the dependence of the tuned hadronization model parameters on variations of $Q_{0}$. As already mentioned in the introduction, the least we need to demand from an improved $S_{mn}$ is a smooth change across multiplicities in the sense that higher partonic multiplicities at lower $Q_{0}$ lead to comparable effect as smaller multiplicities at higher $Q_{0}$ do: the dynamics governing $S_{mn}$ near the infrared cutoff need to be the same as that of the shower close to the infrared cutoff. This puts severe constraints on the onset of the hadronization model, to be discussed in Sec. 4 and Sec. 5. Using the evolution equations implied for $S_{mn}$ for the construction of an evolving model is subject of additional ongoing work.